OpenStudy (anonymous):

Help finding the complex zeros of a function! F(x)=x^4+2x^3+46x^2+98x-147

myininaya (myininaya):

147=3*7*7 --------------- 147=21(7) 147=3(49) Try using synthetic division with the following values x=+ or - {1,3,7,21,49,147}

OpenStudy (anonymous):

Ok, so does it matter what I chose to divide by?

myininaya (myininaya):

It doesn't matter where you start but if you don't get the remainder zero you can throw the number you tried out.

OpenStudy (anonymous):

Perfect!!! Thanks for your assistance you rock!!!

myininaya (myininaya):

Np. I like to at least try to guess ones that are more possible than others and try those because that is a lot to try. lol

OpenStudy (anonymous):

Ok, I did them all, and found there were no complex zeros do you think that's correct!

myininaya (myininaya):

Nope. :(

OpenStudy (anonymous):

Oh man, I was looking at the wrong numbers, I'm redoing my math now :-) one second and ill send what I found :-)

OpenStudy (anonymous):

Okay, -3, and 1 :-)